# Optimized mechanical quadrature squeezing beyond the 3 dB limit via gradient-descent algorithm

### v2

@@ Abstract
The preparation of mechanical quadrature-squeezed states holds significant importance in cavity optomechanics because the squeezed states have extensive applications in understanding fundamental quantum mechanics and exploiting modern quantum technonogy. Here, we propose a reliable scheme for generating mechanical quadrature squeezing in a typical cavity optomechanical system via seeking for optimal cavity-field driving pulses using the gradient-descent algorithm. We realize strong quadrature squeezing in a mechanical resonator that exceeds the $3$ dB steady-state limit, even with a thermal phonon occupancy of one hundred. Furthermore, the mechanical squeezing can be ultrafastly created within one mechanical oscillation period. We also obtain the optimal pulsed drivings associated with the created mechanical squeezings and analyze the mechanism for mechanical squeezing generation. This work will promote the application of optimal quantum control in quantum optics and quantum information science.
@@ I INTRODUCTION
Cavity optomechanics Kippenberg2014RMP is a frontier research field focusing on the radiation-pressure interaction between cavity fields and moving mirrors at macroscopic scales. Owing to the nonlinear properties of the optomechanical interaction and the development of laser cooling techniques Wilson-Rae2007PRL ; Marquardt2007PRL ; ChanNature2011 ; TeufelNature2011 , optomechanical systems provide new paths and opportunities for generation and manipulation of macroscopic nonclassical states, such as entangled states VitaliPRL2007 ; Palomaki2013Science ; Riedinger2016Nature ; Jiao2020PRL ; Yu2020Nature ; Lai2022PRL , quantum superposed states Bose1997PRA ; Marshall2003PRL ; LiaoPRL2016 ; LiaoPRA2020 , and squeezed states of fields FabrePRA1994 ; ManciniPRA1994 ; Brooks2012 ; PurdyPRx2013 ; Safavi-Naeini2013 ; AgarwalNJP2014 and mechanical resonators WoolleyPRA2008 ; HuangPRA2023 ; TeufelNN2009 ; NunnenkampPRA2010 ; LiaoPRA2011 ; MilburnOL1981 ; MariPRL2009 ; PontinPRL2014 ; PontinPRL2016 ; ChowdhuryPRL2020 ; YongChun2023PRA ; ClerkNJP2008 ; SzorkovszkyPRL2011 ; VasilakisNP2015 ; MengPRL2020 ; PirkkalainenPRL2015 ; GenesPRA2009 ; AgarwalPRA2016 ; AsjadPRA2014 ; LuPRA2015 ; RashidPRL2016 ; KronwaldPRA2013 ; XiongOL2018 .
@@
Quantum squeezing, as an essential quantum resource, plays an increasingly significant role in modern quantum technology ranging from quantum precision measurement to quantum information processing Braunstein2003 ; Drummond2004 . Quantum squeezing of mechanical modes holds significance as it has the potential to enhance the accuracy of quantum measurements CavesRMP1980 . There exist many theoretical and experimental schemes for generating mechanical squeezing based on various methods, such as parametric squeezing NunnenkampPRA2010 ; LiaoPRA2011 ; MilburnOL1981 ; MariPRL2009 ; PontinPRL2014 ; PontinPRL2016 ; ChowdhuryPRL2020 ; YongChun2023PRA , quantum measurement ClerkNJP2008 ; SzorkovszkyPRL2011 ; VasilakisNP2015 ; MengPRL2020 ; PirkkalainenPRL2015 , quantum state transfer GenesPRA2009 ; AgarwalPRA2016 ; HuangPRA2023 , mechanical nonlinearities NunnenkampPRA2010 ; AsjadPRA2014 ; LuPRA2015 , and reservoir engineering KronwaldPRA2013 ; XiongOL2018 . Although great advances have been made in squeezing generation, how to generate mechanical squeezing beyond the 3 dB limit AgarwalPRA2016 ; MilburnOL1981 ; Scully1997 in a typical optomechanical system is still an important and desired task in this field. In addition, from the view point of realistic application, an ultrafast generation of strong mechanical squeezing remains a challenge.
@@
In cavity optomechanical systems, the external optical driving provides an effective way to control the quantum properties and dynamical behaviors of the system VannerPNAS2011 ; VannerNC2013 ; LiaoPRA2011 ; MariPRL2009 ; MeenehanPRX2015 . The design of an optimal optical driving to achieve a supposed goal is an interesting topic in this system. Currently, quantum control techniques have been successfully applied to various schemes in cavity optomechanical systems Hammerer2017AIMO , such as robust state transfer Wang2021AP , optimized cooling Jacobs2011 ; MagriniNature2021 ; LiuPRA2022 , entanglement enhancement Stefanatos2017QST ; ClarkeNJP2020 , and strong squeezing Xiong2020PR ; Basilewitsch2019AQE . Furthermore, optimal control technology based on the gradient-descent algorithm Dong2020 has been recognized as a powerful method to accomplish complex control tasks, and hence it is expected to provide new ways to optimize the control of cavity optomechanical systems.
@@
In this paper, we apply the gradient-descent algorithm to prepare mechanical quadrature squeezing in a typical cavity optomechanical system. Our scheme can break the $3$ dB steady-state squeezing limit even when the environment thermal phonon occupation associated with the mechanical resonator reaches the order of one hundred. The mechanical quadrature squeezing is achieved by designing a proper pulsed driving to the cavity field. Concretely, we use the gradient-descent algorithm to iteratively optimize the pulsed driving for achieving strong mechanical squeezing. The optimal waveforms of the pulsed driving amplitude and phase are also obtained. In particular, mechanical squeezing can be ultrafastly prepared within one mechanical oscillation period, providing more chance for realistic applications of the generated squeezing before decoherence. Our method will encourage further researchs on optimal quantum control in cavity optomechanical systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the Hamiltonians and present the equations of motion. In Sec. III, we show the generation of mechanical quadrature squeezing using the gradient-descent algorithm. In Sec. IV, we present some discussions on the physical mechanism and experimental implementation of this scheme and conclude this work. An Appendix is presented to show the derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ used in the gradient-descent algorithm.
@@ II HAMILTONIANS AND EQUATIONS OF MOTION
where $\rho$ is the density matrix of the optomechanical system, $\mathcal{D}[o]\rho=o\rho o^{{\dagger}}-(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)/2$ (for $o=a$, $a^{{\dagger}}$, $b$, and $b^{{\dagger}}$) is the Lindblad superoperator Agarwal2013 , and $H_{\mathrm{opt}}(t)$ is defined in Eq. (II). The parameters $\kappa$ and $\gamma$ are, respectively, the damping rates of the cavity field and the mechanical oscillation, and $\bar{n}_{\mathrm{m}}$ is the environmental thermal-excitation occupation of the mechanical resonator.
@@
where $\rho^{\prime}(t)$ represents the density operator in the displaced representation, $\alpha(t)$ and $\beta(t)$ are the displacement amplitudes of the cavity field and the mechanical resonator, respectively. By substituting $\rho(t)=D_{b}^{\dagger}(\beta)D_{a}^{\dagger}(\alpha)\rho^{\prime}(t)D_{a}(\alpha)D_{b}(\beta)$ into Eq. (2) and setting the coefficients of the driving terms to be zero, we obtain the quantum master equation in the displaced representation as
@@
We point out that both the coefficient matrix $\mathbf{M}(t)$ and the inhomogeneous term $\mathbf{N}(t)$ are functions of the displacement amplitudes $\alpha(t)$ and $\beta(t)$, then the dynamic evolution of the second-order moments can be controlled via adjusting the pulsed driving.
Algorithm. Mechanical squeezing generation
@@ III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
For squeezing-generation task, our goal is to reduce the variance $\Delta X_{b}^{2}(\theta,t)$ such that it is smaller than 1/2. The control strategy of squeezing generation is summarized in $\mathbf{Algorithm}$. Note that a variable learning rate method is adopted here to reduce the iteration number and to improve the efficiency. The detailed calculation of $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ is shown in Appendix.
Figure 2: (a) The loss function $\Delta X_{b}^{2}(\theta=\pi/2,T)$ as a function of the iteration number. The squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ under different target squeezing degrees: (b) $\mathcal{S}_{b}=1$ and (c) $\mathcal{S}_{b}=3.2$. (d,e) The Wigner functions of the generated states for the mechanical mode corresponding to the maximal squeezing in panels (b,c). (f,g) The scaled driving amplitude $\Omega(t)/\omega_{\mathrm{m}}$, (h,i) phase $\phi(t)/2\pi$, and (j,k) mean phonon number $\langle b^{\dagger}b\rangle$ as functions of the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to panels (b,c). Other parameters used are the same as those given in Fig. 1.
The mechanical quadrature squeezing can be well quantized by the squeezing degree AgarwalPRA2016
@@
We display in Fig. 2(a) the loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number to verify the efficiency of the gradient-descent algorithm. Here, we can see that the loss function $\Delta X_{b}^{2}(\pi/2,T)$ decreases gradually as the iteration number increases. The transient mechanical squeezing degrees $1$ dB and $3.2$ dB are obtained at the iteration numbers $1185$ and $4000$, respectively. The corresponding dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ are shown in Figs. 2(b) and 2(c). It shows that there are no mechanical squeezing for a long duration of time firstly, and then oscillatingly increase to the target value in a short period of time. In particular, the squeezing degree $\mathcal{S}_{b}=3.2$ breaks the $3$ dB steady-state limit [Fig. 2(c)], and stronger squeezings can be realized as the iteration number increases. Note that we consider the transient squeezing here rather than steady-state squeezing. Therefore, our results are not conflicting with the $3$ dB steady-state squeezing limit, which is determined by the dynamic stability.
To confirm the mechanical squeezing in phase space, we introduce the Wigner function of the mechanical mode, defined as Vogel2006
@@
where $\mathbf{D}=(D_{\mathrm{R}},D_{\mathrm{I}})^{\mathrm{T}}$ represents the two-dimensional vector, and $\mathbf{V}_{b}$ is the covariance matrix for the mechanical mode. The Wigner functions corresponding to 1 dB and 3.2 dB squeezing are shown in Figs. 2(d) and 2(e). We can see that the squeezing appears along the $\pi/2$-axis (corresponding to the squeezing angle $\theta=\pi/2$), and the larger the squeezing degree, the stronger the quadrature squeezing.
In Figs. 2(f) and 2(g), we show the time-dependent driving amplitudes, which are required to achieve the squeezing degrees $\mathcal{S}_{b}=1$ dB and 3.2 dB, respectively check . We can see that a larger driving amplitude is required to realize a stronger squeezing. In addition, for a larger mechanical squeezing degree, the driving phase oscillation becomes more intense, as shown in Figs. 2(h) and 2(i). We also investigate the dependence of the mean phonon number $\langle b^{\dagger}b\rangle$ (in the displaced representation, associated with the quantum fluctuation) as a function of the scaled time $\omega_{\mathrm{m}}t$ in Figs. 2(j) and 2(k). We observe that $\langle b^{\dagger}b\rangle$ decreases from one hundred to less than one for both $\mathcal{S}_{b}=1$ dB and 3.2 dB. Note that these results can be well understood by examining the statistical properties for the squeezed vacuum state. For a single-mode squeezed vacuum state, the mean phonon number is $\sinh^{2}r$, which is a function of the squeezing amplitude $r$. Further, the covariance of the mechanical rotating-quadrature operator in the squeezing-vacuum state is $(2\sinh^{2}r+1-2\sinh r\cosh r\cos\varphi)/2$, ($\varphi$ is the difference between the rotation angle and the squeezing angle). Considering $\varphi=0$, then we can numerically verify that the squeezing amplitude corresponding to the squeezing degree $\mathcal{S}_{b}=3$ is $r=0.347$, and the mean phonon number is 0.125. This indicates that the mechanical squeezing exists even if the mechanical resonator is cooled near the ground state.
@@
To explore the influence of the deviations in $\Omega(t)$ and $\phi(t)$ on mechanical quadrature squeezing, we introduce the relative deviation $\eta$, defined as $\eta=(Q_{r}-Q_{t})/Q_{t}$. Here $Q$ represents either $\Omega$ or $\phi$, $Q_{r}$ and $Q_{t}$ stand for the realistically used parameters and the learned theoretical parameters, respectively. In Fig. 3(a), we plot the squeezing degree $\mathcal{S}_{b}$ as a function of $\eta$. Here, we see that for the relative deviation $\eta\in[-0.1,0.1]$, the $\mathcal{S}_{b}$ is an increasing function of $\eta$ for both the driving amplitude and phase. In addition, the squeezing is more sensitive to amplitude deviation than phase deviation. In Fig. 3(b), we further investigate the influence of the environment on the squeezing generation once the pulsed driving is removed at $T=120\omega_{\mathrm{m}}^{-1}$. The $\mathcal{S}_{b}$ will experience a slightly decrease from 3.2 to 2.51 during the period from $\omega_{\mathrm{m}}t=120$ to 300, indicating that the squeezing has a good robustness.
@@
Figure 4: (a) The loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number under $T=6\omega_{\mathrm{m}}^{-1}$. The inset shows the local magnification of the loss function over the iteration interval [600,6180]. (b) Evolution of the squeezing degree $\mathcal{S}_{b}$ in one mechanical oscillation period. The inset shows the evolution of $\mathcal{S}_{b}$ over time duration [5.9895,6]. (c) The driving amplitude $\Omega(t)/\omega_{\mathrm{m}}$ and (d) phase $\phi(t)/2\pi$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ after the last iteration ($m=6180$) under $T=6\omega_{\mathrm{m}}^{-1}$. (e) The mean phonon number $\langle b^{\dagger}b\rangle$ as a function of the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to panel (b). Other parameters used are the same as those given in Fig. 1.
Ultrafast generation of mechanical squeezing is a desired task from the viewpoint of transient evolution. Below, we investigate the generation of mechanical squeezing within one mechanical oscillation period. Figure 4(a) shows the loss function $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ versus the iteration number related to $\mathcal{S}_{b}=1$. Here we can see that the $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ gradually decreases from the initial value 319 to 0.397 as the iteration number increases, verifying the validity of the gradient-descent algorithm. The dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ after the last iteration ($m=6180$) is shown in Fig. 4(b). We see that the squeezing occurs in the last extremely short duration due to the harsh time condition, but it is still continuous, as shown in the inset of Fig. 4(b). The results in Figs. 4(c) and 4(d) indicate that a large driving amplitude is required to achieve $\mathcal{S}_{b}=1$ within one mechanical oscillation period and the phase has no distinct signature. We also exhibit in Fig. 4(e) the mean phonon number $\langle b^{\dagger}b\rangle$ versus the scaled evolution time at this time. Here we see that $\langle b^{\dagger}b\rangle$ will be larger than the initial value due to the larger driving amplitude, and then goes to a cooled state with dozens of mean phonons. This indicates that the creation of mechanical squeezing does not necessarily require the ground-state cooling of the mechanical resonator.
We also investigate the influence of the sideband-resolution parameter on the squeezing generation. In Figs. 5(a-c), we plot the squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to different sideband-resolution parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. We show the driving amplitude and phase as functions of $\omega_{\mathrm{m}}t$ for different sideband-resolution parameters in Figs. 5(d-f) and Figs. 5(g-i). To maintain $\mathcal{S}_{b}=1$, a larger driving amplitude is needed for a larger value of $\kappa/\omega_{\mathrm{m}}$, and the phase oscillation becomes more severe for a larger $\kappa/\omega_{\mathrm{m}}$. The corresponding dynamic evolution of the mean phonon number $\langle b^{\dagger}b\rangle$ is shown in Fig. 5(j). Here we can see that the mechanical resonator is cooled from the initial one hundred phonons to a few phonons. The better the sideband-resolution parameter is selected, the deeper cooling of the mechanical resonator. Note that a large $\kappa/\omega_{\mathrm{m}}$ will result in fast photon dissipation within the cavity, thereby affecting the optomechanical interaction and leading to the failure of squeezing generation.

### v3

# Optimized mechanical quadrature squeezing beyond the 3-dB limit via a gradient-descent algorithm

@@ Abstract
The preparation of mechanical quadrature-squeezed states holds significant importance in cavity optomechanics because the squeezed states have extensive applications in understanding fundamental quantum mechanics and exploiting modern quantum technology. Here, we propose a reliable scheme for generating mechanical quadrature squeezing in a typical cavity optomechanical system via seeking optimal cavity-field driving pulses using the gradient-descent algorithm. We realize strong quadrature squeezing in a mechanical resonator that exceeds the $3$-dB steady-state limit, even with a thermal phonon occupancy of 100. Furthermore, the mechanical squeezing can be ultrarapidly created within one mechanical oscillation period. We also obtain the optimal pulsed drivings associated with the created mechanical squeezings and analyze the mechanism for mechanical squeezing generation. This paper will promote the application of optimal quantum control in quantum optics and quantum information science.
@@ I INTRODUCTION
Cavity optomechanics [1] is a frontier research field focusing on the radiation-pressure interaction between cavity fields and moving mirrors at macroscopic scales. Owing to the nonlinear properties of the optomechanical interaction and the development of laser cooling techniques [2, 3, 4, 5], optomechanical systems provide new paths and opportunities for generation and manipulation of macroscopic nonclassical states, such as entangled states [6, 7, 8, 9, 10, 11], quantum superposed states [12, 13, 14, 15], and squeezed states of fields [16, 17, 18, 19, 20, 21] and mechanical resonators [22, 40, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 41, 42, 23, 43, 44].
@@
Quantum squeezing, as an essential quantum resource, plays an increasingly significant role in modern quantum technology ranging from quantum precision measurement to quantum information processing [45, 46]. Quantum squeezing of mechanical modes holds significance as it has the potential to enhance the accuracy of quantum measurements [47]. There exist many theoretical and experimental schemes for generating mechanical squeezing based on various methods, such as parametric squeezing [25, 26, 27, 28, 29, 30, 31, 32], quantum measurement [33, 34, 35, 36, 37], quantum state transfer [38, 39, 40], mechanical nonlinearities [25, 41, 42], and reservoir engineering [43, 44]. In particular, some methods [26, 29, 30, 34, 38, 41, 42, 43] can be used to generate the quadrature squeezing, which exceeds the 3-dB squeezing limit [27, 48]. The strong squeezing has wide applications in quantum technology, and its generation is a significant task in quantum optics. In addition, from the view point of realistic application, an ultrafast generation of strong mechanical squeezing remains a challenge.
@@
In cavity optomechanical systems, the external optical driving provides an effective way to control the quantum properties and dynamical behaviors of the system [49, 50, 26, 28, 51]. The design of an optimal optical driving to achieve a supposed goal is an interesting topic in this system. Currently, quantum control techniques have been successfully applied to various schemes in cavity optomechanical systems [52], such as robust state transfer [53], optimized cooling [54, 55, 56], entanglement enhancement [57, 58], and strong squeezing [60, 59]. Furthermore, optimal control technology based on the gradient-descent algorithm [61] has been recognized as a powerful method to accomplish complex control tasks, and hence it is expected to provide new ways to optimize the control of cavity optomechanical systems.
@@
In this paper, we apply the gradient-descent algorithm to prepare mechanical quadrature squeezing in a typical cavity optomechanical system. Our scheme can break the $3$-dB steady-state squeezing limit even when the environment thermal phonon occupation associated with the mechanical resonator reaches the order of 100. The mechanical quadrature squeezing is achieved by designing a proper pulsed driving to the cavity field. Concretely, we use the gradient-descent algorithm to iteratively optimize the pulsed driving for achieving strong mechanical squeezing. The optimal waveforms of the pulsed driving amplitude and phase are also obtained. In particular, mechanical squeezing can be ultrarapidly prepared within one mechanical oscillation period, providing more chance for realistic applications of the generated squeezing before decoherence. Our method will encourage further researches on optimal quantum control in cavity optomechanical systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the Hamiltonians and present the equations of motion. In Sec. III, we show the generation of mechanical quadrature squeezing using the gradient-descent algorithm. In Sec. IV, we present some discussions concerning the influences of driving amplitude and phase noises on the squeezing generation, the physical mechanism of the present squeezing-generation method, the experimental implementation of this squeezing-generation scheme, and the applications of the optimization methods in cavity optomechanics. Finally, we conclude this paper in Sec. V. The Appendix shows the derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ used in the gradient-descent algorithm.
@@ II HAMILTONIANS AND EQUATIONS OF MOTION
where $\rho$ is the density matrix of the optomechanical system, $\mathcal{D}[o]\rho=o\rho o^{{\dagger}}-(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)/2$ (for $o=a$, $a^{{\dagger}}$, $b$, and $b^{{\dagger}}$) is the Lindblad superoperator [62], and $H_{\mathrm{opt}}(t)$ is defined in Eq. (II). The parameters $\kappa$ and $\gamma$ are, respectively, the damping rates of the cavity field and the mechanical resonator, and $\bar{n}_{\mathrm{m}}$ is the environmental thermal-excitation occupation of the mechanical resonator.
@@
where $\rho^{\prime}(t)$ represents the density operator in the displaced representation, and $\alpha(t)$ and $\beta(t)$ are the displacement amplitudes of the cavity field and the mechanical resonator, respectively. By substituting $\rho(t)=D_{b}^{\dagger}(\beta)D_{a}^{\dagger}(\alpha)\rho^{\prime}(t)D_{a}(\alpha)D_{b}(\beta)$ into Eq. (2) and setting the coefficients of the driving terms to be zero, we obtain the quantum master equation in the displaced representation as
@@
We point out that both the coefficient matrix $\mathbf{M}(t)$ and the inhomogeneous term $\mathbf{N}(t)$ are functions of the displacement amplitudes $\alpha(t)$ and $\beta(t)$, thus the dynamic evolution of the second-order moments can be controlled via adjusting the pulsed driving.
III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
Algorithm1. Mechanical squeezing generation
@@ III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
For the squeezing-generation task, our goal is to reduce the variance $\Delta X_{b}^{2}(\theta,t)$ such that it is smaller than 1/2. The control strategy of squeezing generation is summarized in $\mathbf{Algorithm1}$. Note that a variable learning rate method is adopted here to reduce the iteration number and to improve the efficiency. The detailed calculation of $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ is shown in Appendix.
Figure 2: (a) The loss function $\Delta X_{b}^{2}(\theta=\pi/2,T)$ as a function of the iteration number. The squeezing degree $\mathcal{S}_{b}$ vs the scaled evolution time $\omega_{\mathrm{m}}t$ under different target squeezing degrees: (b) $\mathcal{S}_{b}=1$ and (c) $\mathcal{S}_{b}=3.2$. (d),(e) The Wigner functions of the generated states for the mechanical mode corresponding to the maximal squeezing in panels (b) and (c). (f),(g) The scaled driving amplitude $\Omega(t)/\omega_{\mathrm{m}}$, (h),(i) phase $\phi(t)/2\pi$, and (j),(k) mean phonon number $\langle b^{\dagger}b\rangle$ as functions of the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to panels (b) and (c). Other parameters used are the same as those given in Fig. 1.
The mechanical quadrature squeezing can be well quantized by the squeezing degree [39]
@@
We display in Fig. 2(a) the loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number to verify the efficiency of the gradient-descent algorithm. Here, we can see that the loss function $\Delta X_{b}^{2}(\pi/2,T)$ decreases gradually as the iteration number increases. The transient mechanical squeezing degrees $1$ and $3.2$ dB are obtained at the iteration numbers $1185$ and $4000$, respectively. The corresponding dynamic evolutions of the squeezing degree $\mathcal{S}_{b}$ are shown in Figs. 2(b) and 2(c). It shows that there is no mechanical squeezing for a long duration of time, first, and then oscillation increases to the target value in a short period of time. In particular, the squeezing degree $\mathcal{S}_{b}=3.2$ breaks the $3$-dB steady-state limit [Fig. 2(c)], and stronger squeezings can be realized as the iteration number increases. Note that we consider the transient squeezing here rather than steady-state squeezing. Therefore, our results are not conflicting with the $3$-dB steady-state squeezing limit, which is determined by the dynamic stability.
To confirm the mechanical squeezing in phase space, we introduce the Wigner function of the mechanical mode, defined as [63]
@@
where $\mathbf{D}=(D_{\mathrm{R}},D_{\mathrm{I}})^{\mathrm{T}}$ represents the two-dimensional vector, and $\mathbf{V}_{b}$ is the covariance matrix for the mechanical mode. The Wigner functions corresponding to 1- and 3.2-dB squeezing are shown in Figs. 2(d) and 2(e). We can see that the squeezing appears along the $\pi/2$ axis (corresponding to the squeezing angle $\theta=\pi/2$), and the larger the squeezing degree the stronger the quadrature squeezing.
@@
In Figs. 2(f) and 2(g), we show the time-dependent driving amplitudes, which are required to achieve the squeezing degrees $\mathcal{S}_{b}=1$ and 3.2 dB, respectively [64]. We can see that a larger driving amplitude is required to realize a stronger squeezing. In addition, for a larger mechanical squeezing degree, the driving phase oscillation becomes more intense, as shown in Figs. 2(h) and 2(i). We also investigate the dependence of the mean phonon number $\langle b^{\dagger}b\rangle$ (in the displaced representation, associated with the quantum fluctuation) as a function of the scaled time $\omega_{\mathrm{m}}t$ in Figs. 2(j) and 2(k). We observe that $\langle b^{\dagger}b\rangle$ decreases from 100 to less than 1 for both $\mathcal{S}_{b}=1$ and 3.2 dB. Note that in our simulations, the mechanical resonator is assumed initially in a thermal state at the same temperature with the heat bath. The density matrix for the thermal state is given by $\rho_{\mathrm{th}}=\sum_{n=0}^{\infty}P_{n}|n\rangle\langle n|$, where $P_{n}=\bar{n}_{\mathrm{th}}^{n}/(\bar{n}_{\mathrm{th}}+1)^{n+1}$ represents the probability distribution, with the mean thermal phonon number $\bar{n}_{\mathrm{th}}=\mathrm{Tr}(\rho_{\mathrm{th}}b^{\dagger}b)=1/(e^{\hbar\omega_{\mathrm{m}}/k_{B}T}-1)$.
@@
Figure 4: (a) The loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number under $T=6\omega_{\mathrm{m}}^{-1}$. The inset shows the local magnification of the loss function over the iteration interval [600,6180]. (b) Evolution of the squeezing degree $\mathcal{S}_{b}$ in one mechanical oscillation period. The inset shows the evolution of $\mathcal{S}_{b}$ over time duration [5.9895,6]. (c) The driving amplitude $\Omega(t)/\omega_{\mathrm{m}}$ and (d) phase $\phi(t)/2\pi$ vs the scaled evolution time $\omega_{\mathrm{m}}t$ after the last iteration ($m=6180$) under $T=6\omega_{\mathrm{m}}^{-1}$. (e) The mean phonon number $\langle b^{\dagger}b\rangle$ as a function of the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to panel (b). Other parameters used are the same as those given in Fig. 1.
To explore the influence of the deviations in $\Omega(t)$ and $\phi(t)$ on mechanical quadrature squeezing, we introduce the relative deviation $\eta$, defined as $\eta=(Q_{r}-Q_{t})/Q_{t}$. Here $Q$ represents either $\Omega$ or $\phi$, and $Q_{r}$ and $Q_{t}$ stand for the realistically used parameters and the learned theoretical parameters, respectively. In Fig. 3(a), we plot the squeezing degree $\mathcal{S}_{b}$ as a function of $\eta$. Here, we see that for the relative deviation $\eta\in[-0.1,0.1]$, the $\mathcal{S}_{b}$ is an increasing function of $\eta$ for both the driving amplitude and phase. In addition, the squeezing is more sensitive to amplitude deviation than phase deviation. In Fig. 3(b), we further investigate the influence of the environment on the squeezing generation once the pulsed driving is removed at $T=120\omega_{\mathrm{m}}^{-1}$. The $\mathcal{S}_{b}$ will experience a slight decrease from 3.2 to 2.51 during the period from $\omega_{\mathrm{m}}t=120$ to 300, indicating that the squeezing has a good robustness. Notably, after the external drive is removed, the photons in the cavity will dissipate completely within a certain period, and the optomechanical interaction will no longer work. Consequently, the mechanical resonator will be thermalized by the heat bath, leading to a reduction of the mechanical quadrature squeezing.